OpenStudy (anonymous):

Question 3 (Multiple Choice Worth 2 points) If we know the force, the distance, and the time we can calculate only power. work and power. only work. neither work or power. -------------------------------------------------------------------------------- Question 4 (Multiple Choice Worth 2 points) If you are given time and distance, you can determine power if you know force. watts. energy. joules.

OpenStudy (anonymous):

i already solved these man

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

well,you know that power = work / time?

OpenStudy (anonymous):

yes so my answer would be b

OpenStudy (anonymous):

yes,your answer is correct :)

OpenStudy (anonymous):

can u help with more

OpenStudy (anonymous):

you mean the second question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok,first of all say your answer... then i'll tell you...

OpenStudy (anonymous):

energy i think

OpenStudy (anonymous):

TIME

OpenStudy (anonymous):

well,let's see...energy = 5j , distance = 5m and the tome is = 5sec too.please determine power.

OpenStudy (anonymous):

I MEANT WATT

OpenStudy (anonymous):

@Nick123cross1 ,you shouldn't have SAID the answer if you know!please HELP ...

OpenStudy (anonymous):

ok sorry.

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

Question 5 (Multiple Choice Worth 2 points) In order for work to take place the energy present must be related to the movement of the object. the force applied must be a balanced force resulting in motion of the object. the force applied must cause the movement of the object in the same direction as the force. the movement of the object must be the cause of the force. -------------------------------------------------------------------------------- Question 6 (Multiple Choice Worth 2 points) Less work was done. What is one possible reason why this has happened? There was more force applied; therefore, there was less movement. There was less force applied; therefore, there was more movement. There was less force applied; therefore, there was less movement. There was more force applied; therefore, there was more movement. -------------------------------------------------------------------------------- Question 7 (Multiple Choice Worth 2 points) If we increase the distance traveled when doing work , and keep all other factors the same, what will happen? The amount of force will go down. The amount of force will increase. The amount of work will decrease. The amount of work will increase. -------------------------------------------------------------------------------- Question 8 (Multiple Choice Worth 2 points) For work to be accomplished we must have an action which causes movement and an opposite reaction. an action which causes no movement or reaction. a force and a movement in the opposite direction of the force. a force and a movement in the same direction as the force. -------------------------------------------------------------------------------- Question 9 (Multiple Choice Worth 2 points) A man pulls on his dog's leash to keep him from running after a bicycle. Which term best describes this example? negative work positive work negative force positive force -------------------------------------------------------------------------------- Question 10 (Multiple Choice Worth 2 points) Which statement best describes how work and power are different? To find work we need to know force and distance; to find power we need to know force and velocity. To find work we need to know energy and time; to find power we need to know energy and distance. To find work we need to know velocity and distance; to find power we need to know distance and time. To find work we need to know distance and force; to find power we need to know energy and velocity.

OpenStudy (anonymous):

well, you have to answer all of these questions ?! :D

OpenStudy (anonymous):

come on bro i need help im not good at science

OpenStudy (anonymous):

if you exert a force of 700n to walk 6m up a flight of stairs in 6s, how much power do you use

OpenStudy (anonymous):

well,first of all question 10 : what's your idea about it?

OpenStudy (anonymous):

c i think

OpenStudy (anonymous):

if you exert a force of 700n to walk 6m up a flight of stairs in 6s, how much power do you use

OpenStudy (anonymous):

please pay attention: work ===> N x M ==> we just need to know FORCE and DISTANCE,and sometimes we can work these out!like this: we have velocity and time ==> velocity = m/s and tome = s so velocity x time = m = distance! and power ===> watts OR j/s ==> we need to know energy (OR work , as work is j too) and time.just like work, we can find power in other cases , for example we have force and velocity ===> velocity = m/s and force = N so velocity x force = m.N/s = j/s = watts... in this way you can find the answers for other questions...got it?

OpenStudy (anonymous):

now just let me to write all of the formulas for power...i'll attach it.

OpenStudy (anonymous):

woooooooooow

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

Ok, \[P = \frac{ f . d }{ t } = f . v = \frac{ E }{ t } = \frac{ m.g.h }{ t } \]

OpenStudy (anonymous):

m.g.h and 1/2 mv^2 is energy and usually you won't need them...but some times you should find p in this way... got it?

OpenStudy (anonymous):

anything else?

OpenStudy (anonymous):

@Swagged.Out.Jay

OpenStudy (anonymous):

nope thnks